OpenStudy (fools101):

Have one more and that all! What is the slope of a segment with endpoints (4, 3) and (-2, -2)? A) 2 B) 1/2 C) 6/5 D) 5/6

OpenStudy (fools101):

please help!

OpenStudy (kira_yamato):

\[G = \frac{3 - (-2)}{4 - (-2)} = 5/6\] D

OpenStudy (fools101):

Thank Kira Yamato ^ ^!!! : p